### v1

# Thermoelectric energy conversion in molecular junctions out of equilibrium

Abstract
Understanding time-resolved quantum transport is crucial for developing next-generation quantum technologies, particularly in nano- and molecular junctions subjected to time-dependent perturbations. Traditional steady-state approaches to quantum transport are not designed to capture the transient dynamics necessary for controlling electronic behavior at ultrafast time scales. In this work, we present a non-equilibrium Green’s function formalism, within the recently-developed iterated generalized Kadanoff-Baym ansatz ($i$GKBA), to study thermoelectric quantum transport beyond the wide-band limit approximation (WBLA). We employ the Meir-Wingreen formula for both charge and energy currents and analyze the transition from Lorentzian line-width functions to the WBLA, identifying unphysical divergences in the latter. Our results highlight the importance of finite-bandwidth effects and demonstrate the efficiency of the $i$GKBA approach in modeling time-resolved thermoelectric transport, also providing benchmark comparisons against the full Kadanoff-Baym theory. We exemplify the developed theory in the calculation of time-resolved thermopower and thermoelectric energy conversion efficiency in a cyclobutadiene molecular junction.
I Introduction
While thermoelectric devices are traditionally macroscopic and valued for their stable, time-independent performance, a new generation of applications demands precise control of charge and heat transport at atomic and molecular scales, as well as ultrafast time scales. In these regimes, conventional steady-state descriptions break down, and the interplay between energy conversion and quantum coherence must be addressed. Recent studies have explored these challenges in a variety of systems, including ferromagnetic-superconducting interfaces [1], condensed-matter platforms [2], and qubit-based thermal devices [3]. Coherence effects in mesoscopic systems enable applications such as precision thermometry and efficient refrigeration [4], while thermoelectric phenomena at the atomic and molecular scale highlight the interplay between quantum heat flow, electronic transport, and non-equilibrium thermodynamics [5]. Thermoelectric properties of single-molecule junctions have been explored both experimentally and theoretically [6], with a focus on inelastic transport effects and the Seebeck coefficient [7], and the optimization of the thermoelectric figure of merit [8, 9]. Thermoelectric efficiency and energy harvesting have been investigated in multi-terminal quantum-dot systems operating as thermal engines [10, 11, 12, 13, 14, 15]. These advances underline the importance of thermoelectric transport in enabling quantum technologies, energy-efficient devices, and thermal control techniques at the nanoscale.
The study of time-resolved quantum transport is essential for understanding the behavior of electrons in nano- and molecular junctions under external driving mechanisms, such as time-varying voltages, electromagnetic fields, or sudden changes in system parameters [16, 17, 18, 19]. With miniaturization to atomic and molecular scales, quantum devices push their operating frequencies into the terahertz range, reaching electronic time scales on the order of femtoseconds [20, 21]. Hybrid optoelectronic platforms built from low-dimensional materials (e.g., graphene and topological insulators) exploit high carrier mobility and strong light-matter coupling to function in this regime [22] and have demonstrated active control of electromagnetic signals at sub-picosecond scales using current fabrication techniques [23]. This branch of research is especially relevant for the design of next-generation quantum materials and technologies, where controlling electronic behavior at ultrafast time scales is critical [24, 25, 26, 27, 28]. Traditional approaches to quantum transport, relying on linear response theory and steady-state assumptions [29, 30, 31], are not designed to capture the complex dynamics of systems under strong and time-dependent perturbations. Alternative approaches such as renormalization group and quantum Monte Carlo methods are typically perturbative in the system-lead coupling and become computationally prohibitive in higher dimensions [32]. Time-dependent density-functional theory relies on approximate exchange-correlation functionals, which often fail to capture non-local correlations [33, 34, 35], while scattering wave function methods are generally limited to weakly interacting systems [36].
The non-equilibrium Green’s function (NEGF) formalism serves as a powerful framework to study time-dependent processes [37, 38, 39]. By resolving the time evolution of many-body quantum systems via the Kadanoff-Baym equations (KBE), the NEGF formalism allows for tracking the transient behavior of currents, charge distributions, and other observables in response to sudden changes [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. This time-domain analysis is particularly useful for understanding the immediate response of quantum systems to external drives, as well as for studying relaxation dynamics as the system evolves toward the stationary state [51]. Within the NEGF theory, a compelling technique for time-resolved transport is the generalized Kadanoff-Baym ansatz (GKBA) [52, 53], especially in its recently-developed time-linear framework [54, 55, 56, 57, 58, 59]. While the previous works have established and applied a time-dependent Meir-Wingreen formula for the charge current [59, 60], the corresponding formula for the energy current and thermal transport has only very recently been discussed within the time-linear GKBA formulation [61].
The coupling between a quantum system, such as a molecular junction or quantum dot, and its surrounding leads or electrodes (see Fig. 1) is commonly described by a tunneling rate or line-width function. It characterizes how the states in the central region (e.g., the molecular junction) are broadened due to their interaction with the continuum of states in the leads and determines the rate at which electrons tunnel between the quantum system and the leads. The bandstructure or density of states (DOS) of the leads is decisive in shaping the energy dependence of the line-width function. Due to a considerable computational simplification, a common approach is the so-called wide-band limit approximation (WBLA) where the lead DOS is assumed flat or featureless [62, 63]. However, in a more realistic transport setup, the line-width function deviates from the WBLA form, i.e. having some energy dependence, directly influencing the overall transport characteristics.
Figure 1: Molecular junction model considered in this work. Central system (a cyclobutadiene molecule) is coupled to two leads ($\alpha=L,R$) via the frequency-dependent tunneling rates $\Gamma_{\alpha}$, Eq. (16), characterized by the energy centroids $\epsilon_{\alpha}$ and bandwidths $\Omega_{\alpha}$. The leads are held at different temperatures $\beta_{\alpha}$, and external time-dependent voltages $V_{\alpha}(t)$ are applied.
Therefore, treating time dependence, interactions, and energy-dependent lead coupling is crucial for the accurate modeling of thermoelectric quantum transport in driven nanoscale systems. In this work, we develop a non-equilibrium Green’s function framework, expanding on our recent Letter [61], to address these topics on an equal footing. We present a detailed derivation of the iterated generalized Kadanoff-Baym ansatz ($i$GKBA) concentrating on a Lorentzian tunneling-rate function. We employ the Meir-Wingreen formula for both charge and energy currents and analyze the reduction from the Lorentzian to the WBLA description. The conventional GKBA approach, a subset of the $i$GKBA scheme, is found to be insufficient for studying energy currents, as it leads to an unphysical divergence within the WBLA. This issue is resolved by the finite-bandwidth $i$GKBA as demonstrated by extensive benchmark results and comparison with the full Kadanoff-Baym equation approach in a carbon-based molecular junction. Ultimately, we put the developed $i$GKBA scheme to the test by investigating the time-dependent thermoelectric energy conversion in a molecular junction under a temperature gradient.
II Quantum transport setup and observables
II.1 The Hamiltonian
We consider a generic quantum-transport model, consisting of a central system coupled to metallic leads shown in Fig. 1, described by the time-dependent Hamiltonian
$\hat{H}(t)=\sum_{mn}h_{mn}(t)\hat{c}_{m}^{\dagger}\hat{c}_{n}+\frac{1}{2}\sum_{mnpq}v_{mnpq}(t)\hat{c}_{m}^{\dagger}\hat{c}^{\dagger}_{n}\hat{c}_{p}\hat{c}_{q}\\ +\sum_{k\alpha}E_{k\alpha}(t)\hat{c}_{k\alpha}^{\dagger}\hat{c}_{k\alpha}+\sum_{mk\alpha}\left(T_{mk\alpha}(t)\hat{c}_{m}^{\dagger}\hat{c}_{k\alpha}+\mathrm{h.c.}\right),$ (1)
where $\hat{c}_{m}^{\dagger}/\hat{c}_{n}$ create/annihilate electrons at the central region (orbital/spin basis $m,n$), and $\hat{c}_{k\alpha}^{\dagger}/\hat{c}_{k\alpha}$ create/annihilate electrons at leads (momentum/spin state $k$ at lead $\alpha$). Here, $h_{mn}$ and $v_{mnpq}$ are the one- and two-electron matrix elements of the central region, $E_{k\alpha}$ is the energy dispersion of the leads, and $T_{mk\alpha}$ the tunneling matrix elements between the central region and leads.
The time-dependence of Eq. (1) describes an out-of-equilibrium transport process generated by, e.g., a voltage or coupling switch. The lead energies are modified by a time-dependent bias voltage as $E_{k\alpha}(t)=E_{k\alpha}+V_{\alpha}(t)$. Similarly, a time-dependent (possibly non-local) potential profile within the central region is modeled by $h_{mn}(t)=h_{mn}+u_{mn}(t)$. In addition, we allow for the coupling and interaction matrix elements, $T_{mk\alpha}(t)=T_{mk\alpha}s_{\alpha}(t)$ and $v_{mnpq}(t)=v_{mnpq}s_{v}(t)$ to also have a time-parameter to describe, e.g., adiabatic switching through some time-dependent ramp functions $s(t)$.
II.2 Non-equilibrium Green’s function theory
The central object of the non-equilibrium Green’s function theory, in describing the out-of-equilibrium quantum-transport setting of Eq. (1) is the one-electron lesser ($<$) Green’s function (GF) [38]
$G_{rs}^{<}(t,t^{\prime})=i\left\langle\hat{c}_{s}^{\dagger}(t^{\prime})\hat{c}_{r}(t)\right\rangle,$ (2)
[…]
The GF obeys the integro-differential Kadanoff-Baym equation (in matrix form)
$\left[i\partial_{t}-h^{\mathrm{HF}}(t)\right]G^{<}(t,t^{\prime})$
[…]
where $h_{mn}^{\mathrm{HF}}(t)=h_{mn}(t)+V_{mn}^{\mathrm{HF}}(t)$ is the one-electron Hamiltonian including the Hartree-Fock (HF) potential $V_{mn}^{\mathrm{HF}}(t)=\sum_{pq}(v_{mpqn}(t)-v_{mpnq}(t))\rho_{qp}(t)$. The self-energy $\Sigma=\Sigma_{c}+\Sigma_{\mathrm{em}}$ is composed of the correlation and embedding parts. The correlation self-energy can be specified at different levels of approximations: second-order Born [64, 65], $GW$ and $T$-matrix [66, 67], or based on the non-equilibrium dynamical mean-field theory [68, 69, 70]. The embedding self-energy accounts for the system being open to an environment consisting of the leads. The superscripts ‘$\mathrm{A}$’ and ‘$\mathrm{R}$’ denote the advanced and retarded components, respectively.
Extracting the equal-time limit of Eq. (II.2) gives the equation of motion for the density matrix [59]
$i\frac{d}{dt}\rho(t)=\left[h^{\mathrm{HF}}(t)\rho(t)-iI_{c}(t)-iI_{\mathrm{em}}(t)\right]-\mathrm{h.c.},$ (4)
where $I_{c}$ and $I_{\mathrm{em}}$ are the collision integrals (convolutions) of the correlation and embedding self-energies, respectively, with the GF of the central region. Analogously, interactions of electrons with phonons and photons can be incorporated in the formalism leading to a set of coupled equations for the electron and boson correlators [71, 72, 73].
The equation of motion for the density matrix in Eq. (4) is not closed because the collision integrals depend functionally on the two-time Green’s functions and self-energies. In Sec. III, we use the reconstruction equation [52] to access the two-time lesser GF in terms of simpler propagators, which allows us to develop a time-linear scaling framework.
II.3 Charge and energy currents
We define a current operator for lead $\alpha$ satisfying the Heisenberg equation of motion
$\hat{J}_{\alpha}^{\nu}(t)=\frac{d\hat{H}_{\alpha}^{\nu}}{dt}=i\left[\hat{H},\hat{H}_{\alpha}^{\nu}\right],$ (5)
where $\hat{H}_{\alpha}^{\nu}\equiv\sum_{k}E_{k\alpha}^{\nu}\hat{c}_{k\alpha}^{\dagger}\hat{c}_{k\alpha}$ corresponds to the $\alpha$-lead part of the Hamiltonian (1) when $\nu=1$ and to the particle number operator $\hat{N}_{\alpha}\equiv\sum_{k}\hat{c}_{k\alpha}^{\dagger}\hat{c}_{k\alpha}$ when $\nu=0$. The commutator has non-zero entries only with the coupling block of the full Hamiltonian (1) and we obtain for the currents, $J_{\alpha}^{\nu}\equiv\langle\hat{J}_{\alpha}^{\nu}\rangle$,
$J_{\alpha}^{\nu}(t)=\sum_{mk}E_{k\alpha}^{\nu}\left[T_{mk\alpha}(t)i\bigl{\langle}\hat{c}_{m}^{\dagger}\hat{c}_{k\alpha}\bigr{\rangle}\right.\\ \left.-T_{k\alpha m}^{*}(t)i\bigl{\langle}\hat{c}_{k\alpha}^{\dagger}\hat{c}_{m}\bigr{\rangle}\right],$ (6)
where the operators are understood in the Heisenberg picture. Here, we identify the equal-time lesser GF of the lead-molecule block, cf. Eq. (2): $G_{k\alpha m}^{<}(t,t)=-\left(G_{mk\alpha}^{<}(t,t)\right)^{*}=i\bigl{\langle}\hat{c}^{\dagger}_{m}(t)\hat{c}_{k\alpha}(t)\bigr{\rangle}$. For symmetric and real tunneling matrices, we may rewrite Eq. (6) as [74]
$J_{\alpha}^{\nu}(t)=2\sum_{mk}E_{k\alpha}^{\nu}{\rm Re}\left[T_{mk\alpha}(t)G_{k\alpha m}^{<}(t,t)\right].$ (7)
By using the equations of motion of the lead-molecule Green’s function with Langreth rules we obtain the Meir-Wingreen formula for the charge and energy current (see Ref. [51])
$J_{\alpha}^{\nu}(t)=2{\rm Re}\text{Tr}\int_{-\infty}^{\infty}dt^{\prime}\left[\Sigma_{\alpha}^{\nu,<}(t,t^{\prime})G^{\mathrm{A}}(t^{\prime},t)\right.\\ \left.+\Sigma_{\alpha}^{\nu,\mathrm{R}}(t,t^{\prime})G^{<}(t^{\prime},t)\right]\equiv 2{\rm Re}\text{Tr}I_{\alpha}^{\nu}(t),$ (8)
where the lead self-energies $\Sigma_{\alpha}^{\nu}$ will be defined in detail in Sec. III. For convenience, we also denote $J_{\alpha}(t)\equiv J_{\alpha}^{0}(t)$ as the charge current and $J_{\alpha}^{\mathrm{E}}(t)\equiv J_{\alpha}^{1}(t)$ as the energy current. The associated heat current is then found as $J^{H}_{\alpha}(t)=J^{E}_{\alpha}(t)-\mu_{\alpha}J_{\alpha}(t)$, where $\mu_{\alpha}$ is the chemical potential of lead $\alpha$.
III Time-linear formulation with Lorentzian leads
III.1 Collision integrals
The correlation self-energy terms appearing in the collision integrals [see Eq. (4)] can be separated and dealt with independently for various many-body effects [54, 55, 56]. Focusing now on the embedding self-energy, it accounts for all processes in which an electron transitions from orbital $m$ of the central molecule to an energy level $k$ in lead $\alpha$, and subsequently returns to orbital $n$ of the molecule. It is thus specified in terms of the lead and coupling Hamiltonians, by a summation over all leads, $\Sigma_{\mathrm{em}}=\sum_{\alpha}\Sigma_{\alpha}^{0}$ with $\Sigma_{\alpha,mn}^{0}(t,t^{\prime})=\sum_{k}T_{mk\alpha}(t)g_{k\alpha}(t,t^{\prime})T_{k\alpha n}(t^{\prime})$, where the lead Green’s function $g$ is taken to be non-interacting  [^1] The retarded and lesser components of the lead GF satisfy the equations of motion: $\left[i\partial_{t}-E_{k\alpha}(t)\right]g^{\mathrm{R}}(t,t^{\prime})=\delta(t-t^{\prime})$, $\left[i\partial_{t}-E_{k\alpha}(t)\right]g^{<}(t,t^{\prime})=0$.. Similarly for the collision integral appearing in Eq. (4), $I_{\mathrm{em}}(t)=\sum_{\alpha}I^{0}_{\alpha}(t)$.
Let us now consider the collision integrals defined by Eq. (8)
[…]
We will utilize symbols $x,y,z$ for intermediate times. Dealing with the lesser GF component in Eq. (9) is a fundamental problem of the NEGF formalism. Our goal is to derive a set of coupled equations of motion (EOM) for Eq. (9), to be co-evolved with Eq. (4), constituting a time-linear scaling framework.
III.2 Reconstruction equations
It is convenient to split the lesser GF into the retarded- and advanced-like ($\mathrm{R}$/$\mathrm{A}$) components
$G^{<}(t,t^{\prime})$ $=\theta(t-t^{\prime})G^{<}_{\mathrm{R}}(t,t^{\prime})-\theta(t^{\prime}-t)G^{<}_{\mathrm{A}}(t,t^{\prime}).$ (10)
For these components, we will use of the reconstruction equations derived by Lipavský, Špička, and Velický [52]:
[…]
and
[…]
One obtaines the “standard” GKBA equations by neglecting the memory corrections (given by the integral over $y$). The iterated GKBA ($i$GKBA) is obtained by iterating the equations further, i.e., inserting the GKBA expression for $G^{<}_{\mathrm{A}}$ from the second equation into the first, and $G^{<}_{\mathrm{R}}$ from the first equation into the second. In order to efficiently deal with different kinds of convolution integrals that appear in this work, we introduce short-hand notations:
[…]
$\bigl{[}af\cdot b\bigr{]}_{\mathrm{R}/\mathrm{A}}(t,t^{\prime})$ $=\theta(\pm(t-t^{\prime}))\bigl{[}af\cdot b\bigr{]}(t,t^{\prime}).$ (14)
This allows to rewrite the $i$GKBA equations as follows:
$G^{<}_{\mathrm{R}}(t,x)$ $=-G^{\mathrm{R}}(t,x)\rho(x)$
$\,+\Bigl{[}G^{\mathrm{R}}\cdot\bigl{[}\Sigma^{<}\cdot G^{\mathrm{A}}+\Sigma^{\mathrm{R}}\rho\cdot G^{\mathrm{A}}\bigr{]}_{\mathrm{R}}\Bigr{]}(t,x),$ (15a)
$G^{<}_{\mathrm{A}}(x,t)$ $=-\rho(x)G^{\mathrm{A}}(x,t)$
$\,-\Bigl{[}\bigl{[}G^{\mathrm{R}}\cdot\Sigma^{<}-G^{\mathrm{R}}\rho\cdot\Sigma^{\mathrm{A}}\bigr{]}_{\mathrm{A}}\cdot G^{\mathrm{A}}\Bigr{]}(x,t).$ (15b)
Equations (15) constitute the first important ingredient of our formalism enabling us to evaluate the embedding collision integral.
III.3 Embedding self-energy components
The embedding self-energy is the second important ingredient of our theory. It encodes microscopic properties of the leads (numbered by the index $\alpha$) and is based on the form of the coupling $T_{mk\alpha}$ between the state $m$ of central region to the state $k$ of the lead $\alpha$ with energy $E_{k\alpha}$. From a mathematical point of view, the Lorentzian line-width function [76]
$\Gamma_{\alpha,mn}(\omega)=2\pi\sum_{k}T_{mk\alpha}\delta(E_{k\alpha}-\omega)T_{k\alpha n}\\ =\gamma_{\alpha,mn}\frac{\Omega_{\alpha}^{2}}{(\omega-\epsilon_{\alpha})^{2}+\Omega_{\alpha}^{2}},$ (16)
offers a number of technical advantages. In the equation above, $\epsilon_{\alpha}$ and $\Omega_{\alpha}$ are the energy centroid and bandwidth, respectively, whereas $\gamma_{\alpha}$ is a constant matrix describing the coupling configuration. The wide-band limit approximation is obtained when the bandwidth approaches infinity $\Omega_{\alpha}\to\infty$, in which case the line-width function becomes frequency independent $\Gamma_{\alpha}(\omega)\approx\gamma_{\alpha}$. Since the tunneling matrix elements acquire time-dependence $T_{mk\alpha}s_{\alpha}(t)$ via the ramp functions $s_{\alpha}(t)$ for lead $\alpha$, a time-dependent prefactor appears
$s_{\alpha}(t)e^{-i\phi_{\alpha}(t,t^{\prime})}s_{\alpha}(t^{\prime})$ $=s_{\alpha}(t)u_{\alpha}(t,t^{\prime})s_{\alpha}(t^{\prime}),$
[…]
$\Sigma_{\alpha}^{\nu,\mathrm{R}}(t,t^{\prime})$ $=\left[\Sigma_{\alpha}^{\nu,\mathrm{A}}(t^{\prime},t)\right]^{\dagger}=-is_{\alpha}(t)u_{\alpha}(t,t^{\prime})s_{\alpha}(t^{\prime})$
×$\qquad\times\mathcal{F}\left[\omega^{\nu}\Gamma_{\alpha}(\omega)\right](t-t^{\prime})\theta(t-t^{\prime}),$ (17a)
$\Sigma_{\alpha}^{\nu,<}(t,t^{\prime})$ $=is_{\alpha}(t)u_{\alpha}(t,t^{\prime})s_{\alpha}(t^{\prime})$
×$\qquad\times\mathcal{F}\left[\omega^{\nu}f_{\alpha}(\omega)\Gamma_{\alpha}(\omega)\right](t-t^{\prime}),$ (17b)
where $f_{\alpha}(\omega)$ is the Fermi-Dirac distribution function
$f_{\alpha}(\omega)$ $=\frac{1}{e^{\beta_{\alpha}(\omega-\mu_{\alpha})}+1},$ (18)
for inverse temperature $\beta_{\alpha}$ and chemical potential $\mu_{\alpha}$.
We will first address the embedding self-energy in the calculation of charge currents and lighten the notation $\Sigma_{\alpha}^{0}=\Sigma_{\alpha}$. Due to the relatively simple form of the line-width function (16), by closing the integration contour in the complex lower, upper half-plane we obtain for $\Sigma_{\alpha}^{\mathrm{R}}(t,x)=s_{\alpha}(t)\bar{\Sigma}_{\alpha}^{\mathrm{R}}(t,x)$, $\Sigma_{\beta}^{\mathrm{A}}(x,t)=\bar{\Sigma}_{\beta}^{\mathrm{A}}(x,t)s_{\beta}(t)$, respectively:
$\bar{\Sigma}_{\alpha}^{\mathrm{R}}(t,x)$ $=-\frac{i}{2}\Omega_{\alpha}\gamma_{\alpha}u_{\alpha}(t,x)s_{\alpha}(x)e^{-i\bar{\epsilon}_{\alpha}(t-x)}\theta(t-x),$ (19a)
$\bar{\Sigma}_{\beta}^{\mathrm{A}}(x,t)$ $=\frac{i}{2}\Omega_{\beta}\gamma_{\beta}e^{i\bar{\epsilon}_{\beta}^{*}(t-x)}s_{\beta}(x)u_{\beta}(x,t)\theta(t-x),$ (19b)
where we introduced $\bar{\epsilon}_{\alpha}=\epsilon_{\alpha}-i\Omega_{\alpha}$, and $1\leq\alpha,\beta\leq N_{\text{leads}}$. By explicit calculation, we find the EOMs:
$i\frac{d}{dt}\bar{\Sigma}_{\alpha}^{\mathrm{R}}(t,x)$ $=\bar{\bar{\epsilon}}_{\alpha}(t)\bar{\Sigma}_{\alpha}^{\mathrm{R}}(t,x)+\frac{\Omega_{\alpha}}{2}\gamma_{\alpha}s_{\alpha}(t)\delta(t-x),$ (20a)
$-i\frac{d}{dt}\bar{\Sigma}_{\beta}^{\mathrm{A}}(x,t)$ $=\bar{\bar{\epsilon}}_{\beta}^{*}(t)\bar{\Sigma}_{\beta}^{\mathrm{A}}(x,t)+\frac{\Omega_{\beta}}{2}\gamma_{\beta}s_{\beta}(t)\delta(x-t),$ (20b)
where $\bar{\bar{\epsilon}}_{\alpha}(t)=V_{\alpha}(t)+\bar{\epsilon}_{\alpha}$, and the equal-time conditions read $\bar{\Sigma}_{\alpha}^{\mathrm{R}}(t,t)=\left[\bar{\Sigma}_{\alpha}^{\mathrm{A}}(t,t)\right]^{\dagger}=-\frac{i}{4}\Omega_{\alpha}\gamma_{\alpha}s_{\alpha}(t)$.
In order to compute the lesser self-energy component we use a standard approach of representing the fermionic distribution function in form of a pole expansion [77]
$f_{\alpha}(\omega)=\frac{1}{2}-\sum_{\ell\geq 1}^{N_{p}}\eta_{\ell}\Bigl{[}\frac{1}{\beta_{\alpha}(\omega-\mu_{\alpha})+i\zeta_{\ell}}\\ +\frac{1}{\beta_{\alpha}(\omega-\mu_{\alpha})-i\zeta_{\ell}}\Bigr{]},\quad\text{with\; $\zeta_{\ell}>0$.}$ (21)
Introducing $\Sigma_{\mathrm{R},\alpha}^{<}(t,x)=s_{\alpha}(t)\bar{\Sigma}_{\mathrm{R},\alpha}^{<}(t,x)$, we obtain after contour integrations the explicit expression:
$\bar{\Sigma}_{\mathrm{R},\alpha}^{<}(t,x)$ $=u_{\alpha}(t,x)s_{\alpha}(x)\sum_{\ell\geq 0}\bar{\eta}_{\alpha\ell}e^{-i\bar{\mu}_{\alpha\ell}(t-x)}$
$\qquad=\sum_{\ell\geq 0}\bar{\eta}_{\alpha\ell}\bar{\bar{\Sigma}}_{\mathrm{R},\alpha\ell}^{<}(t,x),$ (22)
where the expansion coefficients
$\bar{\eta}_{\alpha\ell}$ =$=\begin{cases}i\frac{\gamma_{\alpha}}{2}\Omega_{\alpha}f_{\alpha}(\epsilon_{\alpha}-i\Omega_{\alpha})&\ell=0,\\ -\frac{\eta_{\ell}}{\beta_{\alpha}}\Gamma_{\alpha}\left(\mu_{\alpha}-i\frac{\zeta_{\ell}}{\beta_{\alpha}}\right)&\ell\geq 1;\end{cases}$ (23)
and the exponents $\bar{\mu}_{\alpha\ell}$ are given by
$\bar{\mu}_{\alpha\ell}$ =$=\begin{cases}\epsilon_{\alpha}-i\Omega_{\alpha}&\ell=0,\\ \mu_{\alpha}-i\frac{\zeta_{\ell}}{\beta_{\alpha}}&\ell\geq 1.\end{cases}$ (24)
For the lesser self-energy of advanced type we obtain:
$\bar{\Sigma}_{\mathrm{A},\beta}^{<}(x,t)$ $=u_{\beta}(x,t)s_{\beta}(x)\sum_{\ell\geq 0}\bar{\eta}^{*}_{\beta\ell}e^{i\bar{\mu}^{*}_{\beta\ell}(t-x)}$
$\qquad=\sum_{\ell\geq 0}\bar{\eta}^{*}_{\beta\ell}\bar{\bar{\Sigma}}_{\mathrm{A},\beta\ell}^{<}(x,t).$ (25)
The partial self-energies fulfill the EOMs ($\bar{\bar{\mu}}_{\alpha\ell}(t)=V_{\alpha}(t)+\bar{\mu}_{\alpha\ell}$):
$i\frac{d}{dt}\bar{\bar{\Sigma}}_{\mathrm{R},\alpha\ell}^{<}(t,x)$ $=\bar{\bar{\mu}}_{\alpha\ell}\bar{\bar{\Sigma}}_{\mathrm{R},\alpha\ell}^{<}(t,x),$ (26a)
$-i\frac{d}{dt}\bar{\bar{\Sigma}}_{\mathrm{A},\beta\ell}^{<}(x,t)$ $=\bar{\bar{\mu}}_{\beta\ell}^{*}\bar{\bar{\Sigma}}_{\mathrm{A},\beta\ell}^{<}(x,t),$ (26b)
and the equal-time condition reads $\bar{\bar{\Sigma}}_{\mathrm{R},\alpha\ell}^{<}(t,t)=\bar{\bar{\Sigma}}_{\mathrm{A},\alpha\ell}^{<}(t,t)=s_{\alpha}(t)$.
So far, we dealt with ordinary self-energies, i. e., the $\nu=0$ case. Starting from the Fourier representations (17), the relations between $\nu=0$ and $\nu=1$ self-energies can be established:
$\bar{\Sigma}_{\alpha}^{1,\mathrm{R}}(t,x)$ $=\bar{\epsilon}_{\alpha}\bar{\Sigma}_{\alpha}^{\mathrm{R}}(t,x),$ (27)
$\bar{\Sigma}_{\mathrm{R},\alpha}^{1,<}(t,x)$ $=\sum_{\ell\geq 0}\bar{\mu}_{\alpha\ell}\bar{\eta}_{\alpha\ell}\bar{\bar{\Sigma}}_{\mathrm{R},\alpha\ell}^{<}(t,x).$ (28)
III.4 Working scheme
With all the ingredients, the two cases ($\nu=0,1$) of the collision integral (9) can be written together as:
$I^{\nu}_{\alpha}(t)=s_{\alpha}(t)\Biggl{\{}\sum_{\ell\geq 0}\bar{\mu}_{\alpha\ell}^{\nu}\bar{\eta}_{\alpha\ell}\mathcal{D}_{\alpha\ell}^{c}(t)+\bar{\epsilon}_{\alpha}^{\nu}\mathcal{D}_{\alpha}^{d}(t)\\ +\bar{\epsilon}_{\alpha}^{\nu}\sum_{\beta}\bigl{[}\mathcal{A}_{\alpha\beta}^{a}-\mathcal{A}_{\alpha\beta}^{b}\bigr{]}(t)\Biggr{\}},$ (29)
where new correlators have been introduced: The first term in the collision integral (9) gives rise to
$\mathcal{D}_{\alpha\ell}^{c}(t)$ $=\bigl{[}\bar{\bar{\Sigma}}_{\mathrm{R},\alpha\ell}^{<}\cdot G^{\mathrm{A}}\bigr{]}(t,t).$ (30)
The second collision term in combination with the first term of the reconstruction equation (15b) leads to
$\mathcal{D}_{\alpha}^{d}(t)$ $=\left[\bar{\Sigma}^{\mathrm{R}}_{\alpha}\rho\cdot G^{\mathrm{A}}\right](t,t).$ (31)
Finally, the correction terms in Eq. (15) give rise to
$\mathcal{A}_{\alpha\beta}^{a}(t)$ $=\left[\bar{\Sigma}_{\alpha}^{\mathrm{R}}\cdot\big{[}G^{\mathrm{R}}\cdot\Sigma^{<}_{\beta}\big{]}_{\mathrm{A}}\cdot G^{\mathrm{A}}\right](t,t),$ (32a)
$\mathcal{A}_{\alpha\beta}^{b}(t)$ $=\left[\bar{\Sigma}_{\alpha}^{\mathrm{R}}\cdot\big{[}G^{\mathrm{R}}\rho\cdot\Sigma^{\mathrm{A}}_{\beta}\big{]}_{\mathrm{A}}\cdot G^{\mathrm{A}}\right](t,t).$ (32b)
Figure 2: Derivation of the EOMs for the $i$GKBA correlators. Correlators that originate from the self-energy terms in Eqs. (33) are framed in red. Correlators that appear already at the GKBA level are shaded.
Complementing the EOMs for different self-energy components with the EOMs for retarded
$i\frac{d}{dt}G^{\mathrm{R}}(t,t^{\prime})=\delta(t-t^{\prime})\\ +h^{\mathrm{HF}}(t)G^{\mathrm{R}}(t,t^{\prime})+\bigl{[}\Sigma^{\mathrm{R}}\cdot G^{\mathrm{R}}\bigr{]}(t,t^{\prime}),$ (33a)
and advanced electronic propagators
$-i\frac{d}{dt}G^{\mathrm{A}}(t^{\prime},t)=\delta(t^{\prime}-t)\\ +G^{\mathrm{A}}(t^{\prime},t)h^{\mathrm{HF}}(t)+\bigl{[}G^{\mathrm{A}}\cdot\Sigma^{\mathrm{A}}\bigr{]}(t^{\prime},t),$ (33b)
one can formulate a system of EOMs for the correlators, Fig. 2. As explained in Ref. [61], this requires intermediate correlators $\mathcal{B},\mathcal{C},\mathcal{E}$. We refer to the original work for the full set of 14 coupled equations. It is important to note, however, that these equations simplify considerably if the last convolution terms in Eqs. (33) are neglected. In this case, correlators in red frames are not considered.
IV Benchmark simulations for a quantum-dot system
There are different levels of approximations outlined in Sec. III. Inserting the GKBA iteratively back into the reconstruction equations (11) and (12) resulted in the full set of $14$ EOMs, which is referred to $i$GKBA. Neglecting the correction terms $\mathcal{A}$ in the collision integral (29) reduces to the GKBA framework (only shaded correlators in Fig. 2 are considered) with the embedding self-energies described beyond the WBLA, through the Lorentzian line-width function (16). This finite-bandwidth Lorentzian description can be applied systematically to both the collision integrals (29) and the equations of motion for the advanced propagators (33). We refer to this scheme as “GKBA ($\Omega,\ G^{\mathrm{A}}$)”. Alternatively, the retarded/advanced propagators can be approximated at the wide-band approximation level, i.e., the bandwidth $\Omega$ is finite for the embedding self-energies, but the equations of motion (33) include an effective Hamiltonian $h_{\mathrm{eff}}=h^{\mathrm{HF}}-i\gamma/2$ with the frequency-independent line-width matrix ($\gamma=\sum_{\alpha}\gamma_{\alpha}$; see Eq. (16)). We refer to this as “GKBA ($\Omega,\ h_{\mathrm{eff}}$)” with red-framed correlators in Fig. 2 being additionally neglected. The limiting case, $\Omega\to\infty$ is the wide-band limit approximation, simply referred to as “GKBA (WBLA)”: For the charge density and current, we refer to Ref. [59] and for the energy current, an alternative derivation is presented in the Supplemental Material [78]. The ($i$)GKBA calculations were performed using the cheers code [79, 80].
We focus on a thermoelectric transport setup consisting of a two-terminal quantum dot with a temperature gradient and a time-dependent chemical potential difference. This model, studied analytically in Ref. [81], can capture complex dynamic thermoelectric behavior, and our benchmark calculations illustrate how different operating regimes impact efficiency, charge and energy-current responses of the system, and the validity range of the developed theoretical formulation in Sec. III.
Figure 3: Time-dependent charge density $\rho$ (a), charge current $J_{L}$ (b), and energy current $J_{L}^{E}$ (c) at the left-lead interface with the quantum dot being weakly coupled to the leads, $\gamma=0.1$. Inset in panel (c) displays the equilibrium energy current, before the sudden gate voltage is switched on, in terms of the number of poles.
We first fix the quantum dot energy level to $\epsilon_{0}(t)=0.5+u(t)$ with a sudden gate voltage $u(t)=2.5\theta(t)$. The chemical potentials of the left and right leads are first set to $\mu_{L}=\mu_{R}=0$ and there is no temperature gradient $\beta_{L}=\beta_{R}=10$. In all calculations, we set $\Omega_{\alpha}=\Omega$ for all leads $\alpha$. In Fig. 3, the time-dependent response is analyzed in terms of the charge density on the quantum dot $\rho$ and the charge and energy current at the left-lead interface $J_{L}$, $J_{L}^{E}$, when the quantum dot is weakly coupled to the leads $\gamma_{L}+\gamma_{R}=\gamma=0.1$. We note that $\gamma$ is a scalar for the single-level quantum-dot system. The time evolutions are started from an initially disconnected system $\rho(t\to-\infty)=0$ and adiabatically ramping the coupling between the quantum dot and the leads. As discussed in Ref. [61], with aligned chemical potentials and no temperature gradients, the equilibrium state should exhibit no currents. The comparison of density and charge current to the analytical result of Ref. [81] (“Kara Slimane et al.” in Fig. 3) is excellent already at the WBLA level. When the leads are described with a finite-width Lorentzian, GKBA ($\Omega=10,\ h_{\mathrm{eff}}$) and GKBA ($\Omega=10,\ G^{\mathrm{A}}$), there is a small change in the equilibrium density $\rho(t=0)$. This has no apparent effect on the charge current. The $i$GKBA results follow the same behavior.
For the energy current, Fig. 3(c), the situation changes qualitatively. Although the result of Ref. [81] is obtained by numerical integration, the energy current still converges to a physically expected value at equilibrium $J_{L}^{E}(t=0)=0$. The $i$GKBA calculation matches this fairly well, although the finite-bandwidth $\Omega=10$ does not precisely correspond to the WBLA. In contrast, for the standard GKBA approach, at different levels of approximation, the corresponding result changes with the number of poles in the expansion (21). The main panels in Fig. 3 include $N_{p}=30$ poles, and the pole count is varied at the inset of panel Fig. 3(c), showing a slow divergence (logarithmic) in the case of WBLA. Interestingly, the GKBA calculations with finite-width Lorentzians do converge and the convergence is relatively fast, around $N_{p}\gtrsim 10$ poles, but they converge to an unphysical value ($\neq 0$). The overall mismatch in the energy currents is still fairly minimal at weak coupling.
Figure 4: Same as Fig. 3 but the quantum dot being strongly coupled to the leads, $\gamma=1.0$.
The $i$GKBA scheme accurately resolves the energy currents in the strong coupling regime as well; see Fig. 4. Here, the parameters are otherwise the same as in the weak-coupling calculation (Fig. 3) but the coupling strength is varied so that $\gamma=1.0$. Interestingly, the charge density and current [Fig. 4(a) and Fig. 4(b)] are accurately described by the standard GKBA, but the energy current again diverges for the WBLA case with increasing pole count. We clearly see that when the coupling strength is increased, the standard GKBA at the level of WBLA becomes entirely insufficient. The GKBA calculations with finite-band Lorentzian converge again around $N_{p}\gtrsim 10$ poles, but there is still a noticeable shift in the equilibrium energy current. The $i$GKBA corrects this unphysical behavior, but it is worth noting that the finite-bandwidth Lorentzian with $\Omega=10$ description deviates more from the result of Ref. [81], obtained at the WBLA, when the coupling is stronger. Still, it is required to go beyond GKBA to fix the problem of unphysical energy currents in equilibrium.
Figure 5: Time-dependent energy current $J_{L}^{E}$ at the left-lead interface with the quantum dot being coupled to the leads with $\gamma=0.5$. (a) High temperature $\mu_{L}=\mu_{R}=0$, $\beta_{L}=\beta_{R}=1$; (b) chemical potential drop $\mu_{L}=-\mu_{R}=0.5,\beta_{L}=\beta_{R}=10$; and (c) temperature gradient $\mu_{L}=-\mu_{R}=0.5,\beta_{L}=1,\beta_{R}=10$. Insets display the equilibrium energy current, before the sudden gate voltage is switched on, in terms of the Lorentzian bandwidth $\Omega$.
With charge densities and currents showing excellent agreement across different levels of approximation and with the earlier literature, we turn to the energy-current response in the intermediate coupling regime $\gamma=0.5$, modifying the transport setup accordingly. Fig. 5 displays three different cases (a) aligned chemical potentials $\mu_{L}=\mu_{R}=0$ at high temperature $\beta_{L}=\beta_{R}=1$; (b) chemical potential drop $\mu_{L}=-\mu_{R}=0.5$ at lower temperature $\beta_{L}=\beta_{R}=10$; and (c) chemical potential drop $\mu_{L}=-\mu_{R}=0.5$ with temperature gradient $\beta_{L}=1,\beta_{R}=10$. In all cases, the pole count is set to $N_{p}=40$. The GKBA (WBLA) result is again shifted the most from the equilibrium value. For GKBA, the finite-width Lorentzians capture the correct trend well in all cases, but there is always a shift in the equilibrium value. It is worth pointing out that now the cases in Fig. 5(b) and 5(c) display chemical potential drops and thermal gradients, i.e., the equilibrium energy current being non-zero is physically expected. In all cases, the $i$GKBA approach corrects for the unphysical behavior observed in GKBA. In the insets of Fig. 5 we also show the dependence on the Lorentzian bandwidth $\Omega$. For the equilibrium energy current, the GKBA results are not only unphysical but the convergence toward the WBLA at $\Omega\to\infty$ is fairly slow. In contrast, the $i$GKBA approach accurately resolves the equilibrium energy current, which does not vary significantly with the bandwidth $\Omega$.
Figure 6: Short-time transient behavior of the energy current. The WBLA result displays an abrupt jump at the voltage switch-on time, which can be continuously modeled by applying wider bandwidth Lorentzians. The broadening parameter $\Omega\in[2,100]$ is shown in varying color, from magenta to cyan.
The broadening parameter $\Omega$ in the finite-width Lorentzian has a pronounced impact on the initial transient dynamics, as already evidenced in Figs. 3-5. To further illustrate this behavior, Fig. 6 zooms into the early-time regime of the energy current shown in Fig. 5(a), immediately following the gate-voltage quench at $t=0$. With the analytic results at the WBLA [81], we see that the energy current exhibits an abrupt jump at the time of the voltage switch-on. The jump can be shown to build up continuously by applying a larger broadening $\Omega$ in the $i$GKBA scheme. Similar findings have also been reported in the case of thermomechanical potentials across the junction [63]. At larger times $t\gtrsim 1$, the energy currents calculated via large-bandwidth Lorentzians again resolve the WBLA result accurately.
V Comparison of ($i$)GKBA with the full Kadanoff-Baym equations
We now consider a molecular junction modeled at the Pariser-Parr-Pople [82, 83] (PPP) level, where the kinetic and interaction matrix elements are semi-empirically obtained by fitting to more sophisticated calculations. Specifically, we study a cyclobutadiene molecule attached to donor-acceptor-like leads [84]. This molecule consists of four atomic sites arranged in a ring (Fig. 1) and is modeled using PPP parameters derived from an effective valence shell Hamiltonian [85]. Referring to Eq. (1), the single-particle matrix (in atomic units) is given by
$h=-\begin{pmatrix}0.90286&0.11908&0&0.09772\\ 0.11908&0.90286&0.09772&0\\ 0&0.09772&0.90286&0.11908\\ 0.09772&0&0.11908&0.90286\end{pmatrix},$ (34)
and the two-body interaction is of the form $v_{mnpq}=v_{mn}\delta_{mq}\delta_{np}$, with (in atomic units)
$v=\begin{pmatrix}0.43255&0.20143&0.16515&0.20181\\ 0.20143&0.43048&0.20181&0.16515\\ 0.16515&0.20181&0.43255&0.20143\\ 0.20181&0.16515&0.20143&0.43048\end{pmatrix}.$ (35)
The slight asymmetry in the hopping and interaction matrix elements arises from the fact that cyclobutadiene forms a rectangular rather than a square structure [86]. We note in passing that the physico-chemical properties of the cyclobutadiene molecule are fairly multifaceted [87]. Our aim here is to assess the validity of the $i$GKBA scheme when compared against the Kadanoff-Baym equations, independent of how accurate the PPP model description is, and we turn aside further discussions on, e.g., the antiaromatic nature and the resulting instability of the molecule [88].
The coupling between the molecule and the leads is symmetric: the coupling strength between the molecular sites $1$ and $4$ and the left lead is equal to that between sites $2$ and $3$ and the right lead:
$\gamma_{\alpha,ij}=\delta_{ij}\left[\gamma_{L}^{0}\delta_{\alpha L}(\delta_{i1}+\delta_{i4})+\gamma_{R}^{0}\delta_{\alpha R}(\delta_{i2}+\delta_{i3})\right]$ (36)
with $\gamma_{L}^{0}=\gamma_{R}^{0}\equiv\gamma^{0}/4$. We note that $\gamma_{\alpha}$ is a matrix while $\gamma_{(\alpha)}^{0}$ are scalars indicating the coupling strength. To define energy scales, we use the difference between the highest occupied molecular orbital (HOMO) and the lowest unoccupied molecular orbital (LUMO): $\Delta\equiv E_{\mathrm{LUMO}}-E_{\mathrm{HOMO}}=0.244151$ a.u., calculated at the Hartree-Fock level. For charge neutrality in equilibrium, the chemical potential of the lead-molecule system is set in the middle of the HOMO-LUMO gap: $\mu=-0.118715$ a.u., corresponding to the isolated molecule with two electrons. We consider weak ($\gamma^{0}=\Delta/10$) and strong coupling ($\gamma^{0}=\Delta$), as well as low and high temperatures, defined by $\beta=40/\Delta$ and $\beta=8/\Delta$, respectively. Note that the ‘low’ temperature is still relatively high when converted to physical units, but here, thermal broadening begins to affect the system only when $k_{\mathrm{B}}T=1/\beta$ becomes comparable to the molecular energy levels.
The molecular junction is driven out of equilibrium by a sudden bias voltage on the leads, $V_{\alpha}(t)=V_{\alpha}^{0}\theta(t)$, and we consider weak ($V_{L}^{0}=-V_{R}^{0}=\Delta/8$) and strong ($V_{L}^{0}=-V_{R}^{0}=\Delta/2$) bias voltages. In these situations, the energy centroids of the leads are set to $\epsilon_{L/R}=\mu\pm\Delta/8$ for the weak driving and $\epsilon_{L/R}=\mu\pm\Delta/2$ for the strong driving. The driving parameters are chosen to highlight the capability of the $i$GKBA scheme [61], which does not rely on the uniformity of the lead density of states. The strong driving case is enough to excite a sizable electric current across the HOMO-LUMO gap.
We include the effect of two-body interaction at the Hartree-Fock level, i.e., we neglect the correlation self-energy in Eq. (II.2). This simplified modeling of the molecular junction allows us to explore the new approach with mathematical transparency, facilitating the comparison of the $i$GKBA approach with the full Kadanoff-Baym equations.
Figure 7: Time-dependent charge current at the right-lead interface in the cyclobutadiene molecular junction: (a) weak coupling, low temperature; (b) strong coupling, low temperature; (c) weak coupling, high temperature; (d) strong coupling, high temperature. Solid and dashed lines correspond to the strong and weak driving cases, respectively.
In Fig. 7, we present comparative simulations of the various methods for the time-dependent charge current at the right-lead interface. The benchmark KBE results are based on the method of Ref. [41], which we supplement with the Lorentzian line-width function for the embedding self-energies, in order to consistently compare it with the present development. We take the finite-bandwidth Lorentzians as $\Omega=2\Delta$. In contrast to the adiabatic switching procedure in the ($i$)GKBA methods, the KBE approach determines the coupled equilibrium state by solving the Dyson equation on the imaginary time axis [89, 90, 91]. In practice, the imaginary time grid, $\tau\in[0,-\beta]$, is discretized using a uniform power mesh [92, 93] with $p=6$, $u=8$ (total number of $\tau$-points being $12pu+$ 1) to achieve convergence in the total energy up to the relative error $10^{-6}$ for all the cases considered.
In all the cases reported in Fig. 7, we see that $i$GKBA systematically improves upon the GKBA result, and the agreement with KBE is mostly excellent. While the initial transient response to the sudden bias voltage and the following oscillations are qualitatively similar in all methods, quantitative differences appear. The most prominent improvements of the $i$GKBA are the slow decay toward the stationary state and the value of the stationary current. These are both incorrectly described by the GKBA, especially in the strong-bias cases.
VI Time-resolved thermoelectric energy conversion
Now that the validity range of the $i$GKBA approach has been established, we continue with thermoelectric simulations for the molecular junction of the previous section. In addition to a bias voltage, we apply a temperature gradient with $\beta_{L}\neq\beta_{R}$ over the junction. To address thermoelectric energy conversion, we look at the total charge and heat currents $J^{(H)}_{\mathrm{tot}}=(J^{(H)}_{R}-J^{(H)}_{L})/2$ with $J^{H}_{\alpha}=J^{E}_{\alpha}-\mu_{\alpha}J_{\alpha}$. In the linear-response regime, the relationship between the two is commonly written as [7, 1]
$\begin{pmatrix}J_{\mathrm{tot}}\\ J^{H}_{\mathrm{tot}}\end{pmatrix}=\begin{pmatrix}\mathcal{G}&\mathcal{L}\\ \mathcal{R}&\mathcal{K}\end{pmatrix}\begin{pmatrix}\delta V\\ \delta T\end{pmatrix},$ (37)
where $\mathcal{G}$ and $\mathcal{K}$ are the electrical and thermal conductivities, respectively, and $\mathcal{L}$, $\mathcal{R}$ are the thermoelectric coefficients related to the Seebeck and Peltier effects (connected by the Onsager relations), respectively. Here, $\delta V=V_{L}^{0}-V_{R}^{0}$ and $\delta T=1/\beta_{L}-1/\beta_{R}$. For a more realistic driving protocol, we apply a time-dependent voltage profile, $V_{\alpha}(t)=V_{\alpha}^{0}/[1+\exp(-\kappa t)]$ with $\kappa=5\Delta$, which means that around $t=0$ the bias rapidly yet smoothly increases from 00 to $V_{\alpha}^{0}$.
From Eq. (37) we infer that in the absence of charge current, we may calculate [94]
$S\equiv\frac{\mathcal{L}}{\mathcal{G}}=-\left.\frac{\delta V}{\delta T}\right|_{J_{\mathrm{tot}}=0},$ (38)
which is known as the Seebeck coefficient or thermopower. This enables us to determine time-resolved thermopower by adjusting the bias voltage so that the charge current vanishes. Applying a thermal gradient over the junction, $\delta T>0$, generates a non-zero charge current, which can be ‘countered’ by applying a voltage $\delta V<0$. In practice, this type of thermoelectric energy harvesting involves driving an electric current against a load using a thermal gradient as a driving force [10]. Our motivation here is for dynamically enhanced thermoelectric efficiency in a far-from-equilibrium setting, possibly surpassing steady-state limitations set by static material properties and compatibility conditions [95, 96, 97].
Figure 8: Time-resolved thermoelectricity in the cyclobutadiene molecular junction. (a) Total charge currents through the molecular junction with $\gamma^{0}=\Delta/10$ and $(\beta_{L},\beta_{R})=(8,40)/\Delta$, for varying applied voltages $V\equiv V_{L}^{0}=-V_{R}^{0}$. The inset shows instantaneous current-voltage characteristics indicated by the markers in the main panel; gray markers correspond to calculations not shown in the main panel. A linear fit is employed for obtaining the intercept voltage that suppresses the charge current, indicated by the vertical dotted line. (b) Extracted thermopower for different values of the coupling strength and lead temperatures.
The procedure for finding the time-resolved thermopower is visualized in Fig. 8(a), where the total charge current through the molecular junction is shown with varying voltages. Notably, at the stationary state, there is a non-zero charge current because of the thermal gradient only, without applied voltage. At each instant of time, we collect the charge current values and plot them against the applied voltage, and we observe that these values systematically follow a linear relationship [inset in Fig. 8(a)]. Performing a linear fit for all steps during the time evolution, we can extract the voltage value at which the charge current vanishes. Although the applied voltage must ‘counter’ ($\delta V<0$) the thermal gradient in order to suppress the charge current, we also include positive voltages for a more accurate fit. According to Eq. (38), this procedure then gives us the time-resolved thermopower.
The time-resolved thermopower is shown in Fig. 8(b) for different values of the coupling strength $\gamma^{0}$ and the lead temperatures $(\beta_{L},\beta_{R})$. During the adiabatic switching period ($t<0$), in the absence of applied voltages, the charge currents have saturated to certain non-zero values. The thermopower is then evaluated for times $t>0$ only. It exhibits a rapid drop within a few $\Delta^{-1}$ due to fast charge fluctuations in the molecule. For $\gamma^{0}=\Delta/10$ and $(\beta_{L},\beta_{R})=(8,40)/\Delta$, we see a clear transient enhancement of the thermopower, similar to the one reported in Ref. [94]. The transient behavior thus indicates an increased efficiency of the conversion of thermoelectric energy. Since the parameter space of our thermoelectric transport setup is very large, we do not aim to study exhaustively the optimal conditions for this effect but defer this to future work. In addition, the determined thermopower curves include standard deviations originating from the linear-fit procedure explained above, and we observe that the linear relationship holds very accurately throughout the transient dynamics.
Finally, we study the efficiency of the thermoelectric molecular junction functioning as a heat engine. This can be achieved using the following construction in the linear-response and stationary regimes [see Eq. (37) and Fig. 9(a)]: The thermoelectric molecular junction (hereafter referred to as the system) is connected to an external load and thermalized in the presence of a temperature gradient. Since the system operates under a temperature gradient $\delta T$, it develops an open-circuit Seebeck voltage $V_{o}=-S\delta T$, as extracted in Fig. 8. The system has an internal resistance $R_{s}$ and is connected in series with an external load of resistance $R_{l}$. This configuration forms a closed circuit supporting a total current $J_{\mathrm{tot}}=V_{o}/(R_{s}+R_{l})$.
There are two limiting cases: (1) A short-circuit situation arises during the thermalization when $R_{l}=0$, in which case the voltage at the load is $\delta V=J_{\mathrm{tot}}R_{l}=0$. The corresponding short-circuit current is $J_{s}=V_{o}/R_{s}$. (2) An open-circuit situation occurs when $R_{l}\to\infty$, in which case the voltage at the load equals the Seebeck voltage, $\delta V=V_{o}$, and the current vanishes, as already seen in Fig. 8.
In intermediate cases, the total current takes the form $J_{\mathrm{tot}}=J_{s}-\delta V/R_{s}$ (by Kirchhoff’s junction rule), from which we infer the power output at the load: $J_{\mathrm{tot}}\delta V=(J_{s}-\delta V/R_{s})\delta V$. In the stationary state, the power thus follows a downward-opening parabola in $\delta V$, with a maximum at $\delta V=V_{o}/2$, corresponding to the impedance-matching condition $R_{l}=R_{s}$.
The energy conversion efficiency of the thermoelectric junction is then given by
$\eta(t)=\frac{J_{\mathrm{tot}}(t)\delta V}{J_{L}^{H}(t)},$ (39)
where $J_{L}^{H}$ is the heat current from the left (hot) lead to the molecule. This formulation allows us to evaluate $\eta$ without requiring explicit knowledge of either $R_{s}$ or $R_{l}$. Instead, we parametrize the system in terms of the observable load voltage $\delta V$, which can be held approximately constant even in the transient regime by connecting a capacitor in parallel to the load, and which we scan over the interval $[0,V_{o}]$. Due to the time-dependent nature of the setup, we set $V_{o}$ equal to the mean value of the extracted Seebeck voltages shown in Fig. 8 [78]. This captures both the short- and open-circuit limits, where the efficiency necessarily vanishes due to zero extracted power. The efficiency reaches a maximum at an intermediate value of $\delta V$, allowing a full characterization of the junction performance based solely on measurable quantities.
Figure 9: Thermoelectric energy conversion efficiency in the cyclobutadiene molecular junction. (a) Circuit schematic for the thermoelectric molecular junction connected to an external load; (b,c) time- and voltage-resolved efficiency for coupling strengths $\gamma^{0}=\{\Delta/10,\Delta\}$, respectively.
We focus on two representative cases by fixing the temperature difference as $(\beta_{L},\beta_{R})=(8,40)/\Delta$ and varying the coupling strength $\gamma^{0}=\{\Delta/10,\Delta\}$, for which the extracted mean values of the Seebeck voltages are $V_{o}=\{0.0075,0.0053\}$ a.u., respectively [78]. Using Eq. (39), the time- and voltage-resolved efficiency is shown in Fig. 9(b,c). Toward the stationary state, the efficiency takes a parabolic dependence on the load voltage $\delta V$ as expected. In both coupling cases, the efficiency initially grows linearly with $\delta V$ during the transient but eventually saturates to the parabolic shape. The saturation time is predictably longer in the weak-coupling case, where the thermoelectric efficiency exhibits periodic enhancements over tens of $\Delta^{-1}$. With the same parameters, a similarly enhanced thermopower was observed in Fig. 8(b). For stronger coupling [Fig. 9(c)], the transient oscillations are suppressed, and within $t\sim 10\Delta^{-1}$ the expected parabolic behavior sets in.
Overall, the calculated efficiencies are relatively low ($\lesssim 10$%), which is attributed to the specific molecular structure, coupling configuration, and the applied temperature gradient. For voltages outside the range $[0,V_{o}]$, energy is effectively pumped into the system, and the efficiency drops significantly or becomes negative, as dictated by the direction of currents and the form of Eq. (39). More optimized conditions for efficient thermoelectric energy conversion could be explored within the parameter space of this model, or by considering alternative molecular junction setups, and we defer more thorough investigations of quantum thermodynamics phenomena [98, 99] to future work. Ultimately, the Carnot efficiency $\eta_{C}=1-T_{R}/T_{L}$ sets the fundamental upper limit. This bound is only reached in the idealized limit where the thermoelectric figure of merit $ZT=\mathcal{G}S^{2}T/\mathcal{K}\to\infty$; for any finite $ZT$, the efficiency remains strictly below $\eta_{C}$ [100]. Nonetheless, we observe that the transiently enhanced efficiency can clearly surpass the stationary value.
VII Conclusion
We investigated thermoelectric dynamics in open quantum systems beyond the wide-band limit, specifically focusing on our recently-developed iterated generalized Kadanoff-Baym ansatz [61]. This entailed a time-linear scaling, non-equilibrium Green’s function theory for fast and accurate simulation of open system dynamics, where the environment has a non-flat spectral density, in contrast with the standard wide-band limit approximation. Our approach of iterating the reconstruction equation for the lesser and greater Green’s function includes non-Markovian effects, where the interaction with the environment exhibits memory-dependent behavior.
Applying the GKBA iteratively in the reconstruction equation resulted in a set of 14 coupled equations of motion, referred to as the $i$GKBA scheme. Starting from the full $i$GKBA scheme, different levels of approximation were explored. Neglecting correction terms in the collision integral reduced to the GKBA framework, still incorporating retarded/advanced propagators beyond the wide-band limit. We exemplified the different levels of approximation by focusing on thermoelectric transport in a two-terminal quantum dot device, for which analytical benchmark results were available. The conventional GKBA proved insufficient for energy current calculations due to an unphysical divergence in the WBLA, which was resolved by the finite-bandwidth $i$GKBA scheme. We also found that the short-time transient behavior of the energy current exhibited an abrupt jump, which was correctly resolved by wider-bandwidth Lorentzians in the $i$GKBA scheme.
Finally, we modeled thermoelectric transport in a cyclobutadiene molecular junction — an interacting system — analyzing the non-equilibrium dynamics under both weak and strong coupling, as well as varying temperatures. By comparing with the numerical solution of the full Kadanoff-Baym equations, the $i$GKBA scheme was found to systematically improve upon GKBA by more accurately capturing relaxation dynamics and steady-state currents. Using the transient charge and heat currents, we investigated thermoelectric energy conversion by extracting the time-resolved Seebeck coefficient (thermopower) and the device efficiency. The molecular junction, when operating in the transient regime, exhibited a clear enhancement in thermoelectric conversion efficiency compared to its stationary-state performance.
In summary, our approach advances quantum simulations of time-resolved thermoelectric phenomena, offering accurate and efficient access to the ultrafast dynamics of charge and heat flow at the nanoscale. By overcoming limitations of traditional methods, this bottom-up approach opens the door to predictive modeling of, e.g., transistor technology using two dimensional materials [101, 102] and ultrafast bolometers for qubit readout in quantum computing [103]. Our framework is well-suited to address the complex, time-dependent transport phenomena and energy-efficient device design.
Acknowledgements.
R.T. acknowledges the financial support of the Jane and Aatos Erkko Foundation (Project EffQSim) and the Research Council of Finland through the Finnish Quantum Flagship (Project No. 359240). We also acknowledge CSC–IT Center for Science, Finland, for computational resources.